# Towards Energy-Efficient Distributed Agreement

### v1

@@ 4 Binary Consensus
In this section, we show how to improve over the energy complexity bounds obtained in Section 3, by restricting the domain of the input values to a single bit, while keeping the optimal time complexity of $f+1$ rounds. Recall that, in Algorithm 2, the energy complexity was dominated by the number of committees that each player is a member of, which in turn depends on the size of the individual committees, i.e., $f+1$. Here, we explore the natural strategy of reducing the committee size to achieve a lower energy complexity. While conceptually simple, having smaller committees entails that we can no longer rely on each committee having at least one alive member, requiring a player to spread its value among multiple committees instead. We compensate for this lack of information spreading across committees by making players aware of decisive information, some relay who keep spreading the decisive information to further committees. Assume players may choose 1 over 0 when they have the choice (i.e., they know at least 1 player holds each of those values as their initial value). This situation allows a player to know that everyone may decide 1 when he receives the value, and/or he will crash before the end of the procedure.
@@ 4.1 Description of the Algorithm
We now give a high-level overview of our algorithm; see Algorithm 3 for the detailed pseudo code: During initialization, we use procedure JoinCommittees (see Algorithm 1) to fix $f$ different static committees $C_{1},\ldots,C_{f}$, each of size $\lceil\sqrt{n}\rceil$ in such a way that, if a player $p$ belongs to a committee $C_{i}$, then he does not belong to committees $C_{i+1},\dots,C_{i+\lceil\sqrt{n}\rceil+1}$; here we define $C_{j}=\emptyset$, for all $j>f$, to simplify the notation. A crucial property of being a member of committee $C_{i}$ is that every $p\in C_{i}$ will wake up (at least) in round $i$, and thus we say that round $i$ is the awake round of committee $C_{i}$. Each player has a local variable $Y$ that is initialized with the input bit, in which we store the player’s current decision estimate. Since our algorithm restricts itself to sending messages with bit value $1$, called $1$-valued message, we only ever update $Y$ by setting $Y\leftarrow 1$ if it was 00 before receiving such a message.
Conceptually, the algorithm consists of three phases: Phase 1 consists only of round $1$, in which every player $p$ wakes up and, if $p$ has input value $X=1$, then they send a message with a single bit $1$ to all members of committee $C_{1}$. Any member $q\in C_{1}$ who has $Y=0$ and receives such a message becomes active, in the sense that $q$ remains awake in the subsequent Phase 2 for the next $T=\left\lceil\frac{f+1}{\sqrt{n}}\right\rceil$ rounds.
In Phase 2, we propagate these $1$-value messages across all other committees as follows. s $C_{1},\dots,C_{\lceil\frac{f+1}{\sqrt{n}}\rceil}$ in their respective awake rounds, i.e., in rounds $1,\dots,\lceil\frac{f+1}{\sqrt{n}}\rceil$. If a player $p^{\prime}\in C_{i}$ receives the value $1$ for the first time upon waking up in round $i$ (i.e., $p^{\prime}$ never received $1$ before nor had $1$ as their initial value), then $p^{\prime}$ will become active, and remain awake for the next $\lceil\frac{f+1}{\sqrt{n}}\rceil$ rounds, during which $p^{\prime}$ sends bit $1$ to the next corresponding committees whose members are awake in the respective round.
When reaching round $f+1$ (i.e., Phase 3), all players wake up, and those having $Y=1$ broadcast bit $1$. Finally, every player decides on its current value of $Y$.
@@ 1 Player $p$ wakes up in round 1:
2 Locally compute JoinCommittees($f,\sqrt{n}$)
3 $Y\leftarrow 0$; $T\leftarrow 0$
@@
5      $T\leftarrow\lceil\frac{f+1}{\sqrt{n}}\rceil$
6       Send $1$ to every player in committee $C_{1}$
7if $p$ received value $1$ and $Y=0$ then
8      $Y\leftarrow 1$; $T\leftarrow\lceil\frac{f+1}{\sqrt{n}}\rceil$
9
// Phase 2 (round $2,\ldots,f$)
10 For each round $h\in[2,f]$ do:
@@
12      Player $p$ wakes up in round $h$
@@ 13       $T\leftarrow T-1$
14       Send $1$ to every player in $C_{h}$
15
16if $p\in C_{h}$ then
17      Player $p$ wakes up in round $h$
@@
19            $Y\leftarrow 1$; $T\leftarrow\lceil\frac{f+1}{\sqrt{n}}\rceil$
20
21
// Phase 3 (round f+1)
@@ 22 Player $p$ always wakes up in round f+1
23 if $Y=1$ or $X=1$ then
24      Send $1$ to all
25if received value $1$ or $X=1$ then
26      $Y\leftarrow 1$
27
28Decide on $Y$ at the end of round $f+1$
Algorithm 3 Binary Value Consensus. Code for player $p$.
@@ 26      $Y\leftarrow 1$
Figure 2: Binary Value Consensus execution
@@
Suppose that each player starts with a one-bit input value and that at most $f<n$ players crash. There is a deterministic algorithm that solves consensus in $f+1$ rounds, sends $\mathcal{O}(n^{2})$ messages of constant size. Moreover, the algorithm achieves $\mathcal{O}(\lceil{f}/{\sqrt{n}}\rceil)$ energy complexity.
@@
Every player decides on $Y$ in round $f+1$, hence termination is guaranteed in $f+1$ rounds.
@@
Next, we argue validity: In the case where every player starts with 00, no message is sent throughout the entire execution, causing all players to decide on their input value. Now consider the case where everyone starts with $1$. Every player sets their decision value to their input value $1$ in Line 19, and validity follows.

### v2

@@ 4 Binary Consensus
In this section, we show how to improve over the energy complexity bounds obtained in Section 3, by restricting the domain of the input values to a single bit, while keeping the optimal time complexity of $f+1$ rounds. Recall that, in Algorithm 2, the energy complexity was dominated by the number of committees that each player is a member of, which in turn depends on the size of the individual committees, i.e., $f+1$. Here, we explore the natural strategy of reducing the committee size to achieve a lower energy complexity. While conceptually simple, having smaller committees entails that we can no longer rely on each committee having at least one alive member, requiring a player to spread its value among multiple committees instead. Moreover, it might create an opening for the adversary to strategically crash players, one at each round, such that eventually a round is reached where all the members of a committee (or in general all the players that should receive messages) have already crashed. This allows the adversary to abstain from crashing during that particular round without any player receiving new information, which makes it impossible for the algorithm to terminate after $f+1$ rounds. We compensate for this lack of information spreading across committees by using players aware of decisive information (here, players holding 1) as propagating units for further committees. Assume players may choose 1 over 0 when they have the choice (i.e., they know at least 1 player holds each of those values as their initial value). This situation allows a player to know that everyone may decide 1 when he receives the value, and/or he will crash before the end of the procedure.
@@ 4.1 Description of the Algorithm
We now give a high-level overview of our algorithm; see Algorithm 3 for the detailed pseudo code.
@@
2 $h\leftarrow\min\{f,n^{\prime}-\sqrt{n^{\prime}}+1\}$
3
// $h$ = first round of Phase 3 or 4,
// depending on the relation between
// $f$ and $n^{\prime}$
4 Locally compute first batch of
5 committees for Phases 1 and 2: $C_{1},\cdots,C_{h-1}\leftarrowJoinComm$($h-1,\sqrt{n^{\prime}},n^{\prime}$)
6 Locally compute second batch of
7 committees : $C_{h},\cdots,C_{f}\leftarrowJoinComm$(fh+1,f+1,nf-h+1,f+1,n)
8 $Z\leftarrow 0$; $Y\leftarrow 0$; $T\leftarrow 0$
@@
10  $Y\leftarrow 1$; $T\leftarrow\lceil\frac{f+1}{\sqrt{n^{\prime}}}\rceil$
11  Send 11 to every player in committee $C_{1}$
12if $p$ received value $1$ and $Y=0$ then
13 $Y\leftarrow 1$; $T\leftarrow\lceil\frac{f+1}{\sqrt{n^{\prime}}}\rceil$
// Phase 2 (rounds $2,\ldots,h-1$)
14 For each round $r\in[2,h-1]$ do:
@@
16 player $p$ wakes up in round $r$
17  Send 11 to every player in $C_{r}$
@@ 18  $T\leftarrow T-1$
19
20if $p\in C_{r}$ then
21 player $p$ wakes up in round $r$
@@
23    $Y\leftarrow 1$; $T\leftarrow\lceil\frac{f+1}{\sqrt{n^{\prime}}}\rceil$
24
25
// Phase 3 (rounds $h,\ldots,f-1$)
26 For each round $r\in[n^{\prime}-\sqrt{n^{\prime}}+1,f-1]$ do:
// r = first round of the phase
27 if $Y=1$ and $r=n^{\prime}-\sqrt{n^{\prime}}+1$ then
28  player $p$ wakes up in round $r$
29  Send 1 to every player in $C_{r}$.
30if $T>0$ then
31  player $p$ wakes up in round $r$
32  Send 1 to every player in $C_{r}$.
33  $T\leftarrow T-1$
34if $p\in C_{r}$ then
35  player $p$ wakes up in round $r$
36  if $p$ received value 11 $Z=0$ then
37    $Z\leftarrow 1$ $T\leftarrow 1$
38
39
// Phase 4 (rounds $f$, $f+1$)
// round $f$
40 Player $p$ always wakes up in round $f$
41 if $Y=1$ or $Z=1$ then
42  Send 1 to every player in $C_{f}$
43if $p\in C_{f}$ and $p$ received value 11 then
44  $Y\leftarrow 1$
// round $f+1$
@@ 45 Player $p$ always wakes up in round f+1
46 if $p\in C_{f}$ and $Y=1$ then
47 Send 11 to all
48if received value 11 then
49 Decide 1 at the end of round $f+1$
50else
51  Decide 0 at the end of round $f+1$
52
Algorithm 3 Binary Consensus. Code for player $p$, who has input value $X$.
The algorithm is divided into four phases, and employs committees of players, of size roughly $\sqrt{n}$. During the first phase, players holding 1 as their initial value send it to the first committee. During the second phase, players who received the value 1 become active for $\mathcal{O}(f/\sqrt{n})$ rounds, and propagate this value in each round to the currently awake committee. Depending on the relation between $f$ and $\lfloor\sqrt{n}\rfloor^{2}-\lfloor\sqrt{n}\rfloor+1$, a third phase might happen. In this third phase, players behave the same as in phase 2, except that the committees are of size $f+1$ and the active players propagate the value 1 only if they received it during the last round. The fourth phase consists of two last rounds. In round $f$, every player aware of the value 1 sends it to a committee of size $f+1$, and in the last round, each member of that committee broadcasts this value to all players if it received it on the previous round. This strategy ensures that, assuming an adversary crashes at most $f$ players, if a player decides on 1 in the last round, then there must exist a non-faulty player holding 1 in some earlier round <f+1<f+1. This strategy also ensures a worst case energy complexity of $O(\lceil f/\sqrt{n}\rceil)$. A more detailed description follows next.
During initialization, we use Procedure JoinComm (see Algorithm 1) to fix $f$ different static committees $C_{1},\ldots,C_{f}$. Let $n^{\prime}=\lfloor\sqrt{n}\rfloor^{2}$ be the largest square number smaller or equal to $n$ and let $h=\min\{n^{\prime}-\sqrt{n^{\prime}}+1,f\}$. The size of committees depends on the relation between $f$ and $n^{\prime}$. Specifically, we divide the committees into two parts: $C_{1},\ldots,C_{h-1}$, where each committee is of size $\sqrt{n^{\prime}}$, and $C_{h},\ldots,C_{f}$, where each committee is of size $f+1$. Note that if $f\leq n^{\prime}-\sqrt{n^{\prime}}+1$, then $h=f$ and $C_{h-1}=C_{f-1}$ (in other words, all the committees except for $C_{f}$ are of size $\sqrt{n^{\prime}}$). If $f>n^{\prime}-\sqrt{n^{\prime}}+1$, then there could be more than one committee of size $f+1$. To simplify the notation, we also define $C_{f+1}=P$. A crucial property of being a member of committee $C_{r}$ is that every $p\in C_{r}$ will wake up (at least) in round $r$, and thus we say that round $r$ is the awake round of committee $C_{r}$. Each player has two local variables $Y,Z$ that are initialized with 0, in which we store the player’s current decision estimate (each variable is used in exactly one of two parts in the algorithm explained later). Since our algorithm restricts itself to sending messages with bit value $1$, called $1$-valued message, we only ever update $Y$ and $Z$ by setting $Y\leftarrow 1$ or $Z\leftarrow 1$, if they were 0 before receiving such a message.
Conceptually, the algorithm consists of four phases:
Phase 1 consists only of round $1$, in which every player $p$ wakes up and, if $p$ has input value $X=1$, then it sends a message with a single bit $1$ to all members of committee $C_{1}$. Player $p$ becomes active, in the sense that $p$ remains awake for the next $T=\left\lceil\frac{f+1}{\sqrt{n^{\prime}}}\right\rceil$ rounds, where $T$ is a counter that is decremented by $1$ in each subsequent round. Any member $q\in C_{1}$ who has $Y=0$ and receives such a message becomes active as well.
In Phase 2 (rounds $2,\ldots,h-1$), we propagate these $1$-valued messages across all other committees as follows. If a player $p^{\prime}\in C_{r}$ receives the value $1$ for the first time upon waking up in round $r$ (i.e., $p^{\prime}$ never received $1$ before nor had $1$ as their initial value), then $p^{\prime}$ will become active, and remain awake for the next $\lceil\frac{f+1}{\sqrt{n^{\prime}}}\rceil$ rounds, during which $p^{\prime}$ sends bit $1$ to the next corresponding committees whose members are awake in the respective round. Note that the length of Phase 2 depends on $h$ (or in other words the relation between $f$ and $n^{\prime}$). If $f$ is relatively small ($f\leq n^{\prime}-\sqrt{n^{\prime}}+1$), then $h=f$ and upon completing Phase 2 in round h1=f1h-1=f-1, the players proceed directly to Phase 4 (described below). Otherwise, the players proceed to Phase 3.
In Phase 3 (rounds $h,\ldots,f-1$), the size of the committees are $f+1$ and we propagate $1$-valued message in a similar way to Algorithm 2. During round $h$ (the first round of Phase 3), every player awakes and sends a $1$-valued message to every player in $C_{h}$, if it has $Y=1$. Every player $p\in C_{h}$ that receives a $1$-valued message during round $h$ will awake during round $h+1$ and send a $1$-valued message to $C_{h+1}$. In subsequent rounds of Phase 3, every player $p\in C_{r}$ that receives a $1$-valued message for the first time in Phase 3 (it may have received a $1$-value message in previous phases) during round $r$ will awake in round $r+1$ and send a $1$-valued message to $C_{r+1}$.
In Phase 4 (rounds $f,f+1$), all players wake up. In round $f$, every player that has $Y=1$ or $Z=1$ sends a $1$-valued message to every player in $C_{f}$. Every player in $C_{f}$ that receives a 1-valued message sets $Y=1$. In round $f+1$, a player $p$ broadcast 1 if $p\in C_{f}$ and $Y=1$. A player decides 1 if it receives a $1$-valued message during round $f+1$ and decides 0 otherwise.
Figure 2 schematically illustrates the execution of Algorithm 3.
@@ 51  Decide 0 at the end of round $f+1$
Figure 2: An Example of the Execution of Binary Consensus (Algorithm 3):
Phase 1: Every player with input value $X=1$ sends a message to $p_{i}$, then crashes.
Phase 2: $p_{i}$ is only able to send a message to $p_{j}\in C_{2}$ as it crashes. The same happens in each of the rounds of Phase 2, i.e., the player sending $1$ can only send its message to one member of the next committee.
Phase 3: Assume that $f>n^{\prime}-\sqrt{n^{\prime}}$. All committees are now of size $f+1$. Player $p_{m}$, who received the message in the last round of Phase 2, tries to send a message to everyone in $C_{h}$ but crashes and the message is only received by $p_{n}$; in the remaining rounds of Phase 3, this process continues analogously.
Phase 4: Every alive player with $Y=1$ try to send messages in $C_{f}$ in round $f$, they may crash before fully propagating 1 to $C_{f}$. $p_{a}\in C_{f}$ receives a message. In round $f+1$, $p_{a}$ try to broadcast $1$, and the adversary here can not crash $p_{a}$ without exceeding the threshold $f$. Consensus is reached.
The main result of this section, proved hereafter, is the following.
@@
Suppose that each player starts with a one-bit input value and that at most $f<n$ players crash. There is a deterministic algorithm that solves consensus in $f+1$ rounds, sends $\mathcal{O}(nf)$ messages of constant size. Moreover, the algorithm achieves $\mathcal{O}(\lceil{f}/{\sqrt{n}}\rceil)$ energy complexity.
4.2 Notation and Terminology
Before proving Theorem 4.1, we need to define some crucial terms for the analysis of the algorithm. Recall that $h=\min\{n^{\prime}-\sqrt{n^{\prime}}+1,f\}$. If $f\leq n^{\prime}-\sqrt{n^{\prime}}$, then Phase 3 is not executed, and $h=f$ is the first round of the last phase of the execution (denoted Phase 4 in the algorithm, for rounds $f,f+1$). Otherwise, $h=n^{\prime}-\sqrt{n^{\prime}}+1$ is the round that starts Phase 3.
Consider some round $r$. We say that a player $p$ is triggered in round $r$, if $p$ receives a 1-valued message in round $r$ that causes it to set its timer (and start broadcasting 1 as long as its timer is not 0), or $r\in\{1,h,f+1\}$ and $p$ wakes up in $r$ to send a 1-valued message. The next observation is immediate from the description of the algorithm. {observation} Let $m$ be a specific 1-valued message sent by $p$. There exists a round $r$ such that one of the following is true:
  • $m$ is sent as a consequence of $p$ being triggered in round $r$ by receiving a 1-valued message from some player $p^{\prime}$, in some round following $r$, or
  • $r\in\{1,h,f+1\}$ and $p$ wakes up during round $r$ by the algorithm’s description and sends $m$ (since it has $Z=1$, $Y=1$ or $X=1$, depending on the round).
We refer to the round $r$ guaranteed by Observation 4.2 as the trigger round of $m$. Note that if $r\in\{h,f+1\}$ and $p$ wakes up by the algorithm’s description during $r$ (for having $Y=1$), then $p$ must have been triggered in some prior round $r^{\prime}<r$, in which case the trigger round of $m$ is $r^{\prime}$. We say that $p$ was triggered by $p^{\prime}$ to send $m$ in the former case, as well as in the latter case if $r\in\{h,f+1\}$. We say that $p$ was triggered by the algorithm before round $1$ to send $m$ in the latter case if $r=1$. For convenience, we define $C_{f+1}=V$ (i.e., the set of all players).
{observation}
The following hold:
  1. (1)
    During each round $r\in[h+1,f+1]$, a player sends a 1-valued message only if it was triggered in the previous round, and that can happen at most once.
  2. (2)
    A player can be triggered to send messages during rounds $[1,h]$ only once.
Proof 4.2.
Part (1) follows since a player can only be triggered if $Z=0$, and when it is triggered it sets $Z=1$. Part (2) holds since a player can only be triggered if $Y=0$, upon which $Y$ is set to 1.
@@ 4.2 Notation and Terminology
To prove the correctness of the algorithm, we need to show that it satisfies the validity, termination and agreement requirements. To start the proof, we make the following key observation on the structure of committees. Note that this observation is the reason we use the first $n^{\prime}$ players for the committees of Phase 2 rather then all $n$ players. {observation} $C_{i}=C_{i+\sqrt{n^{\prime}}}$ for every $i\in[1,h-1-\sqrt{n^{\prime}}]$.
Proof 4.3.
This follows from the round robin construction of the committees in Algorithm 1 and the fact that $\sqrt{n^{\prime}}$ is an integer; i.e, we can form $\sqrt{n^{\prime}}$ committees of size $\sqrt{n^{\prime}}$. Then, committee number $\sqrt{n^{\prime}}+1$ will be the same as committee number 1 and so on.
Next, we show that if enough players ($f+1$) receive a $1$-valued message prior to round $f+1$, then agreement on $1$ is guaranteed.
Claim 1.
If a non-faulty player receives a 1-valued message before round $f+1$, then all non-faulty players decide $1$ at the end of the execution.
Proof 4.4.
Let $p$ be a non-faulty player that receives a 1-valued message for the first time during round $r\leq f$. If $r=f$, then $p\in C_{r}$ and during round $f$, $p$ sets $Y\leftarrow 1$. Hence, in round $f+1$, $p$ broadcasts 1 to all players, resulting in every non-faulty processor deciding 1. Otherwise ($r<f$), $p$ awakes in round $f$ and sends 1 to every player in $C_{f}$. Since $|C_{f}|=f+1$, at least one recipient is non-faulty. Following the same reasoning of the previous case, the claim follows.
A direct implication of Claim 1 is the following corollary.
Corollary 4.5.
If $f+1$ players receive a 1-valued message before round $f+1$, then all non-faulty players decide 1 at the end of the execution.
Proof 4.6.
At least one of those $f+1$ players is non-faulty, and thus every non-faulty player decides 1 by Claim 1.
Since we defined a player to be non-faulty if it does not crash before deciding, it suffices to show validity, termination, and agreement for non-faulty players.
@@
Every non-faulty player decides on $Y$ in round $f+1$, hence termination is guaranteed in $f+1$ rounds.
@@
Next, we argue validity: In the case where every player starts with 0, no message is sent throughout the entire execution, causing all players who decide to decide on their input value. Now consider the case where everyone starts with $1$. Since $n\geq f+1$, then by Corollary 4.5, every non-faulty player decides 1, and validity follows.